OpenStudy (anonymous):

sam spends $10 more than half of his money.If this amount of money is the same as $60 more that 1/7 of his money,how much did same have to begin with?

OpenStudy (anonymous):

pleassssssssssssseeeeeeeeeeeeeeeeeeeeeeeeee

OpenStudy (anonymous):

hellllllllllllllllllllllllllppppppppppppppppppp

OpenStudy (anonymous):

Let the total money be = $ x x Then Sam has spent = --- + 10 2 x This is the same as ---- + 60 7

OpenStudy (anonymous):

Equating the two we get x x ---- + 10 = ------ + 60 2 7 x x ---- - ----- = 60 - 10 2 7 7x - 2x ------- = 50 14

OpenStudy (anonymous):

multiplying both sides by 14 (basically to get rid of denominator 14 on LHS), we get 7x - 2x = 50 x 14 so 5x = 700 so x = 700/5 = 140 So, to begin with Sam had $140

OpenStudy (anonymous):

If u r satisfied with the answer, pls click on "good Answer" button to show that question has been answered satisfactorily